OpenStudy (anonymous):

@ganeshie8 @sourwing @jim_thompson5910 @myininaya @mathmale @mathslover @KingGeorge @tHe_FiZiCx99 Can you check ten answers in this thread? :) I would prefer not to just keep opening new ones when I just want my answers checked. :)

OpenStudy (vishweshshrimali5):

Your question ?

OpenStudy (anonymous):

@Zarkon @mathstudent55

OpenStudy (anonymous):

8x = 128 A. x = 16 <--- B. x = 26 C. x = 36 D. x = 120

OpenStudy (eric_d):

correct

OpenStudy (the_fizicx99):

10 questions? o_o

mathslover (mathslover):

Are these test questions?

OpenStudy (anonymous):

Nope School is over for me. I'm doing practice tests all summer. :)

mathslover (mathslover):

Cool. Sorry for doubting. We have to confirm :-)

OpenStudy (anonymous):

s'ok. :)

OpenStudy (anonymous):

Which of the numbers for n makes the following equation TRUE? 3(2n − 1) − 2n = 21 A. 4 B. 5 C. 6 <---- D. 7

OpenStudy (the_fizicx99):

Yes, N = 6 is correct. ._. I just realized I was still in your question in another tab

OpenStudy (anonymous):

If 5b − 25 = 25, what is the value of b? A. There is no solution. B. 0 C. 10 <---- D. 50

OpenStudy (the_fizicx99):

Yes, 10 5b - 25 = 25 +25 +25 5b = 50 /5 /5 b = 10

OpenStudy (anonymous):

Which of the numbers for n makes the following equation TRUE? 4(n + 3) + 7n = 34 A. −2 B. 0 C. 2 <--- D. 4

OpenStudy (the_fizicx99):

yup, N = 2

OpenStudy (the_fizicx99):

You know you can check this for yourself, right? Online calculators like Wolfram also have step by step guides >.>

OpenStudy (anonymous):

Solve the following linear equation for x. 2x + 7 = 2(6 + x) – 4 A. x = -2 B. x = C. infinitely many solutions D. no solution <---

OpenStudy (the_fizicx99):

Correct. 2x + 7 = 12 + 2x - 4 2x + 7 = 8 + 2x When you solve the 2x's cancel out and you can't further solve. So No solutions is correct.

OpenStudy (anonymous):

What are the solutions, if any, of the following equation? 6y − 3 = −6y + 3 A. There is no solution. B. There are an infinite number of solutions. <--- C. 0.5 D. 2

OpenStudy (the_fizicx99):

o_o Actually no, it's not B

OpenStudy (the_fizicx99):

Can you show your attempt to solve for \(\ \sf y \)?

OpenStudy (anonymous):

I wokrd it out with someone else, I know. I just don't know where.

OpenStudy (the_fizicx99):

Hmm, well. Start by moving all of the constants to the left hand side. 6y − 3 = −6y + 3 So add 3 to both sides.  Can you do the same for -6y? It's negative so in order to move it to the right you have to do the inverse which is additions. So add 6y to both sides.